# Stand-In: A Lightweight and Plug-and-Play Identity Control for Video Generation

### v2

@@ Introduction
Plug-and-Play to Other Applications
@@
The proposed conditional image branch is designed based on the LoRA module, which ensures inherent compatibility with other DiT-based models. To validate this, we conducted experiments on the pose-guided video generation task using the VACE framework (Jiang et al. 2025). As illustrated in Figure 8, integrating our method significantly improves the facial identity similarity in the generated videos. This not only demonstrates the plug-and-play nature of our approach but also highlights its robustness in preserving identity consistency.
@@
Our framework is also capable of video face swapping, which can be achieved via zero-shot inpainting (Lugmayr et al. 2022). Figure 9 shows that our method not only achieves high-quality facial identity transfer but also maintains strong temporal consistency across frames, resulting in coherent and high-quality videos.
Ablation Study
Method
Face
  Similarity $\uparrow$
Naturalness $\uparrow$
Prompt
  Following $\uparrow$
w/o RSA 0.022 3.528 19.956
w/o CPM 0.021 3.539 19.901
Full Model 0.724 3.922 20.594
Table 3: Ablation study on the core components of our method. The results show that both the Restricted Self-Attention (RSA) and the Conditional Position Mapping (CPM) are crucial for achieving high performance.
Effectiveness of Restricted Self-Attention
We perform an experiment by removing the LoRA module from the conditional image branch. In this setting, the video latent tokens gather information from the reference image tokens exclusively through a training-free self-attention mechanism. Experimental results in Table 3 demonstrate that this approach fails to preserve identity consistency, confirming the necessity of our restricted self-attention design in the framework.
Effectiveness of Conditional Position Mapping
To validate the effectiveness of the proposed conditional position mapping, we perform an ablation study by removing the reference image’s positional mapping in the 3D RoPE mechanism and retraining the conditional image branch. As shown in Table 3, experimental results demonstrate that eliminating this mapping significantly degrades performance, confirming that conditional position mapping is essential for achieving accurate and robust alignment in our framework.
Conclusion
@@
Supplementary Materials for
Stand-In: A Lightweight and Plug-and-Play Identity Control for Video Generation
Contents
AMore Comparison Results 10
BMore Ablation Study Results 15
CPseudocode 16
DComplete Video Generation Prompts 17
EMore Results 18
Appendix A More Comparison Results
We provide more comparisons of our model with other identity-preserving video generation methods in Figures 11-15.
Figure 11: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (1/5).
Figure 12: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (2/5).
Figure 13: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (3/5).
Figure 14: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (4/5).
Figure 15: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (5/5).
Appendix B More Ablation Study Results
The visual results of our ablation study, presented in Figure 16, together with the quantitative metrics from the main text, demonstrate that both Restricted Self-Attention and Conditional Position Mapping are crucial to our method. A standard, training-free self-attention mechanism is insufficient to establish effective information exchange between the video latents and the conditional image. Furthermore, omitting the positional mapping for the conditional image makes this reference relationship exceedingly difficult for the model to learn.
Figure 16: Visual results of the ablation study.
Appendix C Pseudocode
@@
Appendix D Complete Video Generation Prompts

### v3

@@ 4.3 User Study
4.4 Plug-and-Play to Other Applications
@@
The proposed conditional image branch is designed based on the LoRA module, which ensures inherent compatibility with other DiT-based models. To validate this, we conducted experiments on the pose-guided video generation task using the VACE framework [vace]. As illustrated in Figure 11, integrating our method significantly improves the facial identity similarity in the generated videos. This not only demonstrates the plug-and-play nature of our approach but also highlights its robustness in preserving identity consistency.
@@
Our framework is also capable of video face swapping, which can be achieved via zero-shot inpainting [inpainting]. Figure 12 shows that our method not only achieves high-quality facial identity transfer but also maintains strong temporal consistency across frames, resulting in coherent and high-quality videos.
4.5 Ablation Study
Effectiveness of Restricted Self-Attention (RSA)
To verify the role of RSA, we replaced it with Vanilla Self-Attention (VSA), which concatenates the reference-image and video tokens and allows full bidirectional information flow. Because VSA treats the reference tokens as part of the dynamic video context, the model often fails to refer to the reference image. As shown in Figure 5, this naive design leads to diffused attention and unreliable identity preservation. In contrast, RSA constrains attention so that image queries cannot attend to video tokens. This design keeps the reference representation static while still providing identity cues. As shown in Table 3, RSA improves face similarity from 0.422 to 0.724 and also slightly enhances naturalness, confirming its essential role in maintaining both identity consistency and visual quality.
Table 3: Ablation study on the core components of our method. Replacing Restricted Self-Attention (RSA) with Vanilla Self-Attention (VSA) or Conditional Position Mapping (CPM) with Shared Position Mapping (SPM) degrades both Face Similarity and Naturalness.
Method Face Similarity $\uparrow$ Naturalness $\uparrow$
RSA $\rightarrow$ VSA 0.422 3.855
CPM $\rightarrow$ SPM 0.536 3.755
Full Model 0.724 3.922
Effectiveness of Conditional Position Mapping (CPM)
We further investigate the effect of CPM by Replacing CPM with a Shared Position Mapping (SPM), which places reference and video tokens within the same positional coordinate system. As illustrated in Figure 6, SPM breaks the pretrained positional prior of the backbone model, resulting in unstable spatial layouts. By geometrically separating the coordinate spaces, CPM preserves the pretrained positional prior while maintaining the reference as a global identity prior. In Table 3, CPM increases face similarity from 0.536 to 0.724 and improves naturalness from 3.755 to 3.922, demonstrating its effectiveness in stabilizing video synthesis and enhancing identity fidelity.
5 Conclusion
@@
Stand-In: A Lightweight and Plug-and-Play Identity Control
for Video Generation
     Supplemental Material
6 Pseudocode
@@
7 More Comparison Results
We provide more comparisons of our model with other identity-preserving video generation methods in Figures 14-17.
Figure 14: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (1/4).
Figure 15: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (2/4).
Figure 16: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (3/4).
Figure 17: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (4/4).
8 More Ablation Study Results
To further validate the effectiveness of our proposed Restricted Self-Attention (RSA) and Conditional Position Mapping (CPM), we provide additional qualitative comparisons in Figure 18 . The ablation focuses on two alternative configurations: (1) replacing RSA with vanilla self-attention (VSA), and (2) replacing CPM with Shared Position Mapping (SPM).
When RSA is replaced with Vanilla Self-Attention (VSA), the model suffers from severe identity drift. This confirms that unconstrained bidirectional attention causes the reference-image tokens to be contaminated by dynamic video content, preventing stable identity extraction. Similarly, when CPM is replaced with Shared Position Mapping (CPM → SPM), sharing the positional space results in clearly distorted facial proportions (e.g., enlarged face in the second example), revealing that the pretrained positional prior is severely disrupted. In contrast, our full model consistently maintains both accurate identity and natural facial geometry for both cases, demonstrating that RSA and CPM are jointly crucial for stable, identity-preserving video generation.
Figure 18: Visual results of the ablation study.
9 More Generalization Results
To further evaluate the robustness and generalization ability of our approach, Figure 19 showcases identity-preserving video generation results for a diverse set of subjects. Despite significant variations in age, ethnicity, skin tone, and facial geometry, our method consistently reconstructs key identity traits and maintains stable facial structure across frames. The generated videos exhibit natural motion, coherent expressions, and high perceptual fidelity. The text prompts from top to bottom are:
(1) A white man stands in front of his desk, facing the camera while speaking and gesturing to emphasize his points. He then picks up a pen and writes a few lines on a sheet of paper before looking back up with a confident smile, as if showing his progress to the viewer. The lighting is soft, giving him a natural and self-assured presence.
(2) An elderly Asian woman sits on a wooden balcony chair, smiling gently at the camera. She takes a small sip from her teacup, then looks up with a warm smile and nods. She reaches to adjust the shawl draped over her shoulders and closes her eyes slightly to enjoy the sunlight. Her expression is peaceful and kind, giving a sense of warmth and comfort.
(3) An elderly Black man sits on his home balcony, looking directly at the camera while holding a warm cup of tea with both hands. A gentle breeze brushes past as he softly blows across the tea’s surface and looks up with a warm smile. He then raises his cup and nods toward the camera in a calm and friendly gesture.
(4) A white woman faces the camera, gently sweeping a makeup brush across her face. She then offers a radiant smile, tilting her head slightly to show her makeup before returning to the front and giving a playful wink. Her movements are graceful and natural, complemented by soft lighting.
(5) An Asian man faces the camera, lightly adjusting the collar of his shirt before running his hand through his hair with a smile. He then raises a cup of tea and makes a small “cheers” gesture toward the camera. His expression is natural and confident, and his movements feel relaxed and smooth.
(6) A young Black woman smiles at the camera while adjusting her curly hair. She gently shakes a coffee cup toward the lens, as if sharing her morning energy with the viewer. She blinks playfully, makes an “OK” gesture, and smiles as sunlight from the window falls softly across her face, giving her a confident and natural expression.
(7) A young white boy smiles at the camera, giving a small wave. He glances down with a shy laugh before looking up again with a bright and innocent expression. Sunlight illuminates his face, making him look natural and adorable.
(8) A young Asian girl faces the camera and makes a heart shape with her hands. She then picks up a glass of juice, takes a sip, and smiles as she wipes the corner of her mouth. A breeze lifts her hair as she playfully brushes it aside, her eyes bright and expressive.
(9) A young Black girl smiles brightly at the camera. She waves hello softly, then forms a heart shape with her hands. Her eyes are bright, her smile sweet, and the sunlight behind her creates a warm and gentle atmosphere.
Figure 19: Given a single reference image (left), our method generates videos (right) that maintain strong identity consistency across diverse ethnicities and age groups, despite being trained with only $\sim$1% additional parameters and just 2000 training pairs.
10 Complete Video Generation Prompts